### v2

# RaGS: Unleashing 3D Gaussian Splatting from 4D Radar and Monocular Cues  for 3D Object Detection

Abstract
4D millimeter-wave radar has emerged as a promising sensor for autonomous driving, but effective 3D object detection from both 4D radar and monocular images remains a challenge. Existing fusion approaches typically rely on either instance-based proposals or dense BEV grids, which either lack holistic scene understanding or are limited by rigid grid structures. To address these, we propose RaGS, the first framework to leverage 3D Gaussian Splatting (GS) as representation for fusing 4D radar and monocular cues in 3D object detection. 3D GS naturally suits 3D object detection by modeling the scene as a field of Gaussians, dynamically allocating resources on foreground objects and providing a flexible, resource-efficient solution. RaGS uses a cascaded pipeline to construct and refine the Gaussian field. It starts with the Frustum-based Localization Initiation (FLI), which unprojects foreground pixels to initialize coarse 3D Gaussians positions. Then, the Iterative Multimodal Aggregation (IMA) fuses semantics and geometry, refining the limited Gaussians to the regions of interest. Finally, the Multi-level Gaussian Fusion (MGF) renders the Gaussians into multi-level BEV features for 3D object detection. By dynamically focusing on sparse objects within scenes, RaGS enable object concentrating while offering comprehensive scene perception. Extensive experiments on View-of-Delft, TJ4DRadSet, and OmniHD-Scenes benchmarks demonstrate its state-of-the-art performance. Code will be released.
Introduction
Autonomous driving requires accurate 3D perception for safe decision-making (Mao et al. 2023). Recently, 4D millimeter-wave radar has emerged as a highly promising sensor, known for its robustness in challenging environments, long-range detection capabilities, and ability to capture both velocity and elevation data. Cameras, on the other hand, provide high-resolution semantic information. The complementary nature of these two modalities is crucial for enhancing perception in autonomous driving, particularly for 3D object detecting. However, effective as well as efficient fusion across modalities remains a challenge, as the task is inherently sparse while still requiring a comprehensive understanding of the entire scene.
Figure 1: 4D radar and camera fusion pipelines. (a) Instance-based fusion relies on 2D detection, limiting scene understanding. (b) BEV-based fusion uses predefined grids, causing inefficiencies in background modeling and fixed anchor sampling. (c) Our Gaussian-based fusion offers adaptive sparse objects attention while preserving scene perception.
Figure 2: Pipeline of RaGS. RaGS consists of a Feature Extractor & Head, Frustum-based Localization Initiation (FLI), Iterative Multimodal Aggregation (IMA), and Multi-level Gaussian Fusion (MGF). The positions of the Gaussians are initialized using the FLI module, along with learnable attributes such as rotation, scale, opacity, and implicit feature embeddings. These Gaussians are then passed into the IMA module, where they are projected onto the image plane to gather semantic information. Next, they are processed as voxels using sparse convolution with height-extended radar geometry. Residuals relative to regions of interest are computed iteratively, updating the positions towards sparse objects. Finally, the multi-level Gaussians are rendered into Bird’s Eye View (BEV) features and fused through MGF, followed by cross-modal fusion for 3D object detection.
Recent research has focused on either instance-based or BEV-based fusion for 4D radar and camera 3D object detection. Instance-based approaches (Kim et al. 2023a; Long et al. 2023; Nabati and Qi 2021) rely on 2D detectors to generate proposals, which are refined by aligning radar features. While these approaches directly maintain the 3D detection flow by focusing on sparse objects, they lack global scene understanding and are constrained by cascaded network designs, as shown in Fig. 1(a). On the other hand, BEV-based approaches (Kim et al. 2023b; Xiong et al. 2024; Zheng et al. 2023; Lin et al. 2024; Bai et al. 2024) project multimodal features into a predefined, fixed top-down space, enabling global reasoning. However, these approaches suffer from rigid voxelization, fixed anchor positions for sampling image semantics, and inefficiencies due to excessive background aggregation, as illustrated in Fig. 1(b). While they offer promising performance in building comprehensive scene perception, they are misaligned with the sparse nature of 3D object detection tasks, in which focusing primarily on foreground objects is significantly more efficient.
In parallel, 3D Gaussian Splatting (GS) (Kerbl et al. 2023) has emerged as a compact, continuous scene representation. Originally developed for neural rendering, GS models scenes as collections of anisotropic Gaussians optimized in 3D space, offering physical interpretability, sparsity, and flexibility. Recent extensions have applied GS to dynamic scenes (Yang et al. 2024) and LiDAR-camera fusion (Hess et al. 2025). However, these works have primarily focused on rendering or occupancy prediction (Huang et al. 2024), without exploring its potential for multi-modal fusion or 3D object detection, despite its natural alignment with the flexible needs of such tasks. Unlike 2D instance-based or BEV-based approaches, GS can dynamically allocate resources and adjust its attention, making it particularly efficient for identify interested objects within entire scene.
To address the limitations of existing fusion approaches and better adapt to the sparse nature of 3D object detection, we propose RaGS, a novel framework that leverages 3D Gaussian Splatting to fuse 4D radar and monocular cues, as presented in Fig. 1(c). Unlike traditional dense voxel grids, RaGS represents the scene as a continuous field of 3D Gaussians, allowing for flexible aggregation without the constraints of fixed grid resolutions and fixed anchor sampling. By dynamically allocating resources iteratively, RaGS focuses on sparse foreground objects while preserving global scene perception through Gaussians’s continuous probability distribution. Specifically, we follow a cascaded pipeline composed of three key modules, as presented in Fig. 2. First, the Frustum-based Localization Initiation (FLI) identifies foreground regions and unprojecting pixels to initialize coarse, grounded 3D Gaussian localization. Next, the Iterative Multimodal Aggregation (IMA) aggregates radar and image features within frustum regions, refining Gaussian positions towards objects and enhancing both semantic and geometric expressiveness. Finally, the Multi-level Gaussian Fusion (MGF) integrates features rendered from Gaussians at different levels to produce a rich representation for object prediction. We evaluate RaGS on three 4D radar and camera benchmarks: View-of-Delft (VoD) (Palffy et al. 2022), TJ4DRadSet(Zheng et al. 2022), and OmniHD-Scenes(Zheng et al. 2024), achieving state-of-the-art performance across the datasets, demonstrating the effectiveness of continuous Gaussian representation for multi-modal 3D object detection. Our contributions are summarized as
  • We propose RaGS, the first framework that leverages 3D Gaussian Splatting to fuse 4D radar and monocular cam-era inputs for 3D object detection.
  • We design three modules to construct and refine the Gaussian from multi-modal inputs, enabling dynamic focus on objects while maintaining scene perception.
  • Extensive experiments on View-of-Delft, TJ4DRadSet and OmniHD-Scenes benchmarks demonstrate the effectiveness of our RaGS.
Related Work
4D Radar based 3D Object Detection
Recent work on 3D object detection using 4D radar and camera fusion typically follows either instance-based or BEV-based paradigms. Instance-based approaches generate 2D proposals from images, refined with radar features. Examples include CenterFusion (Nabati and Qi 2021) that aligns radar points with 2D boxes for better center predictions, CRAFT (Kim et al. 2023a) that integrates radar range and velocity cues, and RADIANT (Long et al. 2023) that recovers missed detections under occlusion. While effective for foreground localization, these approaches depend heavily on proposal quality and lack global perception.
BEV-based approaches project multi-modal features into a top-down space for global spatial reasoning. RCFusion (Zheng et al. 2023) utilizes orthographic projection for semantics but neglects depth ambiguity, whereas other approaches like LXL (Xiong et al. 2024) and CRN (Kim et al. 2023b) introduce various improvements for image view transformation. L4DR (Huang et al. 2025b) and V2X-R (Huang et al. 2025a), on the other hand, further explored the role of LiDAR and 4D radar fusion for 3D object detection. Despite their strengths, BEV approaches rely on fixed-resolution voxel grids, leading to inefficiencies, excessive background aggregation, and reduced flexibility in 3D object detection. In contrast, we propose a dynamic 3D Gaussian-based representation, modeling multi-modal features as a continuous Gaussian field. This allows for more flexible and fine-grained fusion, making it particularly suitable for sparse 3D object detection.
3D Gaussian Splatting in Autonomous Driving
3D Gaussian Splatting (GS) has emerged as a key technique for real-time 3D scene reconstruction. Initially introduced as a faster alternative to neural radiance fields for rendering, GS has extended to tasks like non-rigid motion (Yang et al. 2024), large-scale environments (Zhou et al. 2024), and multi-modal fusion with LiDAR-camera (Zhao et al. 2024; Hess et al. 2025) or radar-only inputs (Kung et al. 2025).
Despite these advancements, current GS applications are predominantly in dense prediction tasks, such as rendering or occupancy prediction (Huang et al. 2024). However, its potential for 3D object detection remains largely unexplored, even though the task’s sparse nature inherently aligns with Gaussian flexibility. While NeRF-based approaches (Xu et al. 2023; Tian, Du, and Duan 2023) use volumetric representations for 3D detection, GS offers advantages like faster rendering and better integration with detection pipelines. In this work, we leverage GS for sparse object modeling and scene perception. Unlike traditional BEV grid models that rely on fixed anchors for feature sampling, our method utilizes GS to enable dynamic feature aggregation by adaptively refining Gaussian positions towards targets.
Method
Overview
Fig. 2 illustrates the overall architecture of our RaGS framework, which consists of four main components: Feature Extractor & Head, Frustum-based Localization Initiation (FLI), Iterative Multimodal Aggregation (IMA), and Multi-level Gaussian Fusion (MGF). RaGS encodes image and radar depth into features, and radar point clouds into sparse geometric features. It then initializes 3D Gaussians via FLI, refines them with radar geometry and image semantics via IMA, and finally renders these Gaussians into multi-scale BEV features via MGF for 3D object detection.
Feature Extractor
[…]
Then, the metric depth $\mathbf{D}$ is computed by summing over predefined depth bins $\mathbf{D}=\sum_{d=1}^{D}P_{d}\cdot d$, where $P_{d}$ represents the probability of each depth bin $d$. A lightweight segmentation network is further used to predict foreground regions from image features $\mathbf{F}^{\text{2D}}$, producing segmentation logits $\mathbf{L}\in\mathbb{R}^{H\times W}$. For the 4D radar point cloud, to fully utilize its RCS and velocity characteristics, we pillarize it and apply RCS-Vel-Aware sparse convolution to extract channel-specific information followed (Zheng et al. 2023). These channel-splited features are then stacked into sparse pillars $\mathbf{F}^{\text{pillar}}\in\mathbb{R}^{G\times C}$, where $G$ is the number of pillars. This separate extraction helps avoid confusion and guides the subsequent feature aggregation of the Gaussian.
Gaussian Aggregation
Frustum-based Localization Initiation (FLI). Gaussian initialization is crucial for guiding downstream feature aggregation. Previous approaches use learnable embeddings to initialize localization based on dataset-specific distributions. In this work, we initialize the position of each Gaussian using monocular cues, which provide a coarse but structured foreground prior focused on sparse objects. Thus, the Gaussians are modeled as a combination of two components: explicit physical properties (position $\mathbf{P}$, rotation $\mathbf{R}$, scale $\mathbf{S}$, opacity $\mathbf{O}$) and implicit feature embeddings. The position is inferred from the FLI.
[…]
$\mathbf{P}_{\text{unproj}}=d\cdot\mathbf{K}^{-1}(u,v,1)^{T},$ (1)
where $\mathbf{K}$ is the camera intrinsic matrix. Furthermore, we also include 4D radar points $\mathbf{P}_{\text{radar}}$ and randomly sample points in the frustum space to enrich spatial coverage and eliminate potential instability of foreground identification. For the random sampling process, we begin by defining the candidate points $\mathbf{P}_{\text{cand}}$ as the centers of predefined voxels in the radar coordinate system. These points are projected onto the image using the camera’s extrinsic parameters:
$(ud,vd,d)=\mathbf{K}\left(\mathbf{R}\mathbf{P}_{\text{cand}}+\mathbf{T}\right),$ (2)
where $\mathbf{R}$ and $\mathbf{T}$ are the rotation matrix and translation vector from radar to camera coordinates. We select points that lie within the image boundaries using the condition $\mathbf{I}=\{i\mid 0\leq u_{i}<W,0\leq v_{i}<H\}$, and apply Furthest Sampling:
$\mathbf{P}_{\text{sample}}=\mathtt{FurthestSampling}\left(\mathbf{P}_{\text{cand}}[\mathbf{I}]\right).$ (3)
The final localization $\mathbf{P}$ is a concatenation of the unprojected foreground points, 4D radar points, and sampled points:
$\mathbf{P}=\mathtt{Concat}(\mathbf{P}_{\text{unproj}},\mathbf{P}_{\text{sample}},\mathbf{P}_{\text{radar}})\in\mathbb{R}^{N\times 3}.$ (4)
This ensures that all initialized Gaussians are within the field of view (FoV) and offer coarse-but-structured initial localization, unlike BEV-based approaches that perceive the entire scene. The fully learnable explicit features include a rotation vector $\mathbf{R}\in\mathbb{R}^{N\times 4}$, scale $\mathbf{S}\in\mathbb{R}^{N\times 3}$, and opacity $\mathbf{O}\in\mathbb{R}^{N\times 1}$, which are further concatenated with the position $\mathbf{P}\in\mathbb{R}^{N\times 3}$ to get final explicit physical attributes $\mathbf{F}^{\text{E}}$:
$\mathbf{F}^{\text{E}}=\mathtt{Concat}(\mathbf{P},\mathbf{R},\mathbf{S},\mathbf{O})\in\mathbb{R}^{N\times 11}.$ (5)
Here, $N$ is the number of Gaussian anchors we predefined. Additionally, we assign each Gaussian a learnable embedding $\mathbf{F}^{\text{I}}\in\mathbb{R}^{N\times C}$, which serves as a query for aggregating image semantics and radar geometry. As a result, each Gaussian can be represented $\mathbf{G}=\{\mathbf{\mathbf{F}^{\text{E}},\mathbf{F}^{\text{I}}}\}$.
Figure 3: Procedure of Iterative Multimodal Aggregation (IMA). IMA involves the iterative aggregation of multi-modal features, followed by the updating of Gaussian locations within the frustum.
Figure 4: Dynamic Object Attention of RaGS. We visualize activated Gaussians (approximately 30% of total) in the scene. RaGS focuses on sparse foreground objects while maintaining scene understanding.
Figure 5: Visualization results on the VoD validation set (first row) and TJ4DRadSet test set (second row) . Each figure corresponds to a frame. Orange and yellow boxes represent ground-truths in the perspective and bird’seye views, respectively. Green and blue boxes indicate predicted results. Zoom in for better view.
Figure 6: Qualitative comparison with state-of-the-art 4D radar based 3D object detection approaches.
Iterative Multimodal Aggregation (IMA). To enrich the Gaussians with multi-modal information, we iteratively aggregate features from both radar and image domains, which output $M$ Gaussian set, as presented in Fig. 3. Instead of directly projecting Gaussians into 2D image space and applying 2D deformable attention for semantic aggregation, we first construct a depth-aware 3D image feature space out product image feature $\mathbf{{F}^{\text{2D}}}$ with depth probability $\mathbf{{D}^{\text{prob.}}}$. We then perform 3D deformable cross attention (3D-DCA) within it, enabling the Gaussians to interact with semantically and geometrically aligned image features in a spatially consistent manner. Here we treat each Gaussian as a point located at its mean and conduct projection procedure. More specifically, given 3D Gaussian query $\mathbf{F}^{\text{I}}$ at location P, we aggregate image semantics by
$\mathbf{F}^{\text{I}}=\sum_{n=1}^{T}\mathbf{A}_{n}\mathbf{W}\cdot\phi(\mathbf{{F}^{\text{2D}}}\otimes\mathbf{{D}^{\text{prob.}}},\mathcal{P}(\textbf{P})+\Delta\textbf{q}),$ (6)
where $n$ indexes $T$ sampled offsets around P, $(\textbf{P})$ P ) denotes projection function, $\Delta\textbf{q}\in\mathbb{R}^{3}$ denotes learnable offsets, and $\phi(\cdot)$ performs trilinear interpolation to extract features. The attention weights $\mathbf{A}_{n}\in[0,1]$ and projection matrix $\mathbf{W}$ are learned to guide the aggregation. The utilization of 3D deformable cross-attention (3D-DCA) facilitates coherent interaction between Gaussians in 3D space and image features in the perspective view.
Consequently, we further enhance the implicit feature representation through sparse convolution. Specifically, each Gaussian is treated as a point located at its mean, and the resulting point cloud is voxelized into a sparse grid $\mathbf{V}^{\text{gs}}\in\mathbb{R}^{N\times C}$. Then, instead of relying solely on Gaussian-derived voxels, we additionally incorporate radar pillars $\mathbf{F}^{\text{pillar}}$ to enable synergistic multi-modal learning. Since radar pillars lack vertical resolution, we replicate each pillar along the height dimension to get $\mathbf{V}^{\text{radar}}\in\mathbb{R}^{(G\times Z)\times C}$, here $Z$ is the voxel number along the height dimension, which can be expressed as $\mathbf{V}^{\text{radar}}=\mathtt{Repeat}(\mathbf{F}^{\text{pillar}})$. These radar voxels are inserted into the same voxel grid as the Gaussian-generated ones, and the RCS-velocity aware pillars provide clear physical meaning, thereby better assisting the Gaussians. By concatenating both modalities in a unified voxel space and applying sparse convolution, we treat each gaussian as voxels and enable deep interaction between radar geometry and image semantics. This process is formulated as
$\mathbf{F}^{\text{I}}\leftarrow\mathbf{V}^{\text{gs}}=\mathtt{Spconv}(\mathtt{Concat}(\mathbf{V}^{\text{gs}},\mathbf{V}^{\text{radar}}))[:N],$ (7)
[…]
Specifically, to achieve the finer foreground modeling required for 3D object detection, we refine each Gaussian localization within the frustum, which allows a more precise match with the monocular settings. To achieve this, we first reproject each Gaussian into frustum space and then using an MLP conditioned on its implicit feature $\mathbf{F}^{\text{I}}$ and transformed position $\mathcal{P}(\mathbf{P})$ to predict a residual spatial offset as
$\Delta\mathbf{p}=(\Delta h,\Delta w,\Delta d)=\mathtt{MLP}(\mathtt{Concat}(\mathbf{F}^{\text{I}},\mathcal{P}(\mathbf{P}))),$ (8)
where $\Delta h,\Delta w,\Delta d$ corresponds to vertical, horizontal, and depth adjustments in perspective view and $\mathcal{P}$ denotes the projection from 3D space to the image frustum. The predicted residual is then added to frustum-space location of each Gaussian before being transformed back to 3D space, which ensures consistency with preceding modules. This process is formulated as
$\mathbf{P}\leftarrow\mathcal{P}^{-1}(\mathcal{P}(\mathbf{P})+\Delta\mathbf{p}).$ (9)
This iterative process, depicted in Fig. 4, progressively refines Gaussian localization for sparse objects, in contrast to the fixed grid sampling of previous BEV-based models, which suffer from inflexibility and inefficiency due to the inevitable excessive background aggregation.
[…]
$\mathbf{F}^{\text{gs}}=\mathtt{Conv}(\mathtt{Concat}(\mathbf{F}^{(1)},\ldots,\mathbf{F}^{(L)})).$ (10)
Subsequently, the fused feature is further enhanced with radar sparse pillars $\mathbf{F}^{\text{pillar}}$ through cross-modal fuser to produce final BEV feature that forwarded to the head for detection, which can be expressed as $\mathbf{F}^{\text{BEV}}=\mathtt{CMF}(\mathbf{F}^{\text{gs}},\mathbf{F}^{\text{pillar}})$. Here, $\mathtt{CMF}$ consists of concatenation and convolution layers. During this process, we render the final-layer Gaussians into a depth map, optimize opacity using LiDAR depth, and use the fused BEV features to predict a segmentation map in the bird’s-eye view, supervised by the occupancy map.
Crucially, while MGF rasterizes to BEV, our Gaussians undergo iterative refinement towards object centric. This fundamentally differs from rigid grid by enabling dynamic aggregation that prioritizes sparse foreground objects, thereby avoiding the background inefficiencies inherent to BEV approaches that rely on fixed anchor sampling. The auxiliary losses further regularize this dynamic process (see supplementary material for details).
Approach Input Entire Annotated Area (%) Driving Corridor (%) FPS
Car Ped Cyclist mAP Car Ped Cyclist mAP
ImVoxelNet (WACV 2022) (Rukhovich 2022) C 19.35 5.62 17.53 14.17 49.52 9.68 28.97 29.39 11.1
RadarPillarNet (IEEE TIM 2023) (Zheng et al. 2023) R 39.30 35.10 63.63 46.01 71.65 42.80 83.14 65.86 98.8
SMURF (IEEE TIV 2023) (Liu et al. 2024) R 42.31 39.09 71.50 50.97 71.74 50.54 86.87 69.72 -
SCKD (AAAI 2025) (Xu et al. 2025) R 41.89 43.51 70.83 52.08 77.54 51.06 86.89 71.80 39.3
FUTR3D (CVPR 2023) (Chen et al. 2023) R+C 46.01 35.11 65.98 49.03 76.98 43.10 86.19 69.32 7.3
BEVFusion (ICRA 2023) (Liu et al. 2023) R+C 37.85 40.96 68.95 49.25 70.21 45.86 89.48 68.52 7.1
RCFusion (IEEE TIM 2023) (Zheng et al. 2023) R+C 41.70 38.95 68.31 49.65 71.87 47.50 88.33 69.23 9.0
RCBEVDet (CVPR 2024) (Lin et al. 2024) R+C 40.63 38.86 70.48 49.99 72.48 49.89 87.01 69.80 -
LXL (IEEE TIV 2023) (Xiong et al. 2024) R+C 42.33 49.48 77.12 56.31 72.18 58.30 88.31 72.93 6.1
LXLv2 (IEEE RAL 2025) (Xiong et al. 2025) R+C 47.81 49.30 77.15 58.09 - - - - 6.5
HGSFusion (AAAI 2025) (Gu et al. 2025) R+C 51.67 52.64 72.58 58.96 88.28 62.61 87.49 79.46 -
SGDet3D (IEEE RAL 2025) (Bai et al. 2024) R+C 53.16 49.98 76.11 59.75 81.13 60.91 90.22 77.42 9.2
RaGS (Ours) R+C 58.15 50.81 76.62 61.86 88.15 61.67 88.69 79.50 10.5
Table 1: Comparison with state-of-the-art approaches on the val set of VoD (Palffy et al. 2022). The best values are in bold, and the second-best values are underlined.
Approach Input $\text{AP}_{\text{3D}}$ (%) $\text{AP}_{\text{BEV}}$ (%)
ImVoxelNet C 14.96 17.12
RadarPillarNet R 30.37 39.24
SMURF R 32.99 40.98
FUTR3D R+C 32.42 37.51
BEVFusion R+C 32.71 41.12
LXL R+C 36.32 41.20
LXLv2 R+C 37.32 42.35
RCFusion R+C 33.85 39.76
HGSFusion R+C 37.21 43.23
SGDet3D R+C 41.82 47.16
RaGS (Ours) R+C 41.72 51.04
Table 2: Comparison with state-of-the-art approaches on the test set of TJ4DRadSet (Zheng et al. 2022).
Approach Res. Input mAP(%) ODS(%)
PointPillars - R 23.82 37.20
RadarPillarNet - R 24.88 37.81
Lift, Splat, Shoot 544×960 C 22.44 26.01
BEVFormer 544×960 C 26.49 28.10
PanoOcc 544×960 C 29.17 28.55
BEVFusion 544×960 C+R 33.95 43.00
RCFusion 544×960 C+R 34.88 41.53
RaGS (Ours) 544×960 C+R 35.57 40.94
Table 3: Comparison of 3D object detection on the OmniHD-Scenes (Zheng et al. 2024) test set.
Detection Head and Loss Function
On one hand, following (Bai et al. 2024), we apply the depth loss $\mathcal{L}_{\text{depth}}$ and the perspective segmentation loss $\mathcal{L}_{\text{seg}}$ to supervise the raw feature extraction for pretraining, which can be expressed as
$\mathcal{L}_{\text{pretrain}}=\mathcal{L}_{\text{depth}}+\mathcal{L}_{\text{seg}}.$ (11)
On the other hand, we adopt the 3D object detection loss $\mathcal{L}_{\text{det}}$ following (Zheng et al. 2023), along with auxiliary rendering losses: the rendered depth loss in the perspective view $\mathcal{L}_{\text{depth\_render}}$, and the semantic segmentation loss in the BEV $\mathcal{L}_{\text{seg\_render}}$. The total joint training loss is defined as
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{det}}+\lambda(\mathcal{L}_{\text{depth\_render}}+\mathcal{L}_{\text{seg\_render}}),$ (12)
where $\lambda$ is the hyperparameter balancing detection and auxiliary tasks. In this work, we simply set $\lambda=0.1$. See supplementary material for more details.
Experiments
Implementation Details
Datesets and Evaluation Metrics. We evaluate our model on three 4D radar benchmarks: the VoD (Palffy et al. 2022), TJ4DRadSet (Zheng et al. 2022), and OmniHD-Scenes (Zheng et al. 2024). The VoD dataset, collected in Delft, includes 5,139 training frames, 1,296 validation frames, and 2,247 test frames. The TJ4DRadSet, captured in Suzhou, contains 7,757 frames with 5,717 for training and 2,040 for testing, posing challenges due to complex scenarios such as nighttime and glare. The OmniHD-Scenes dataset features multimodal data from six cameras, six 4D radars, and a 128-beam LiDAR, with 11,921 keyframes annotated, consisting of 8,321 for training and 3,600 for testing.
We follow the official evaluation protocols for each dataset. For VoD, we report 3D Average Precision (AP) across the full area and drivable corridor. For TJ4DRadSet, we evaluate 3D AP and BEV AP within 70 meters of the radar origin. For OmniHD-Scenes, we use mean Average Precision (mAP) and the OmniHD-Scenes Detection Score (ODS) for detection within a ±60 m longitudinal and ±40 m lateral range around the ego vehicle.
Network Settings and Training details. For all datasets, the final voxel is set to a cube of size 0.32 m, and the image sizes are 800×1280 for VoD, 640×800 for TJ4DRadSet, and 544×960 for OmniHD-Scenes. Anchor size and point cloud range are kept as in (Zheng et al. 2024). The models are trained on 4 NVIDIA GeForce RTX 4090 GPUs with a batch size of 4 per GPU. AdamW is used as the optimizer, with 12 epochs for pretraining and 24 epochs for joint training.
3D Object Detection Results
Results on View-of-Delft and TJ4DRadSet Dataset. Table 1 presents the 3D object detection results on the VoD validation set (Palffy et al. 2022), where our RaGS consistently outperforms state-of-the-art 4D radar and camera fusion models. For cars, we achieve exceptional accuracy, leveraging the flexible occupancy capability of the Gaussian model and an effective fusion design. For pedestrians and cyclists, we also deliver near best performance. When compared to the strong BEV-based baseline LXL (Xiong et al. 2024), our method shows significant improvements, with gains of 5.55% in $\text{mAP}_{\text{EAA}}$ and 6.57% in $\text{mAP}_{\text{DC}}$. Furthermore, it achieves a frame rate of 10.5 FPS, demonstrating its suitability for real-world applications. Compared to the VoD dataset, the TJ4DRadSet dataset (Zheng et al. 2022) presents greater challenges, including complex scenarios such as nighttime, under-bridge conditions, and the addition of a truck category with highly variable object sizes. Despite these difficulties, RaGS achieves state-of-the-art performance, with the highest $\text{mAP}_{\text{BEV}}$ of 51.04% and the second-best $\text{mAP}_{\text{3D}}$ of 41.72%, as shown in Table 2. Visualization results of VoD and TJ4DRadSet are shown in the first and second row of Fig. 5, respectively. Qualitative comparison are presented in Fig. 6.
Baseline w/ FLI w/ IMA w/ MGF $\text{mAP}_{\text{EAA}}$ $\text{mAP}_{\text{DC}}$
55.33 72.32
57.40 75.80
59.12 77.68
59.45 76.98
Table 4: Overall ablation of RaGS.
Baseline FLI $\text{mAP}_{\text{EAA}}$ $\text{mAP}_{\text{DC}}$
Frustum Radar Depth
55.33 72.32
56.26 73.10
56.72 75.78
57.40 75.80
Table 5: Ablation study of FLI on the VoD dataset.
FLI IMA $\text{mAP}_{\text{EAA}}$ $\text{mAP}_{\text{DC}}$
3D-DCA Pillars Frustum
57.40 75.80
58.79 76.16
58.93 76.35
59.45 76.98
Table 6: Ablation study of IMA on the VoD dataset.
Results on OmniHD-Scenes Dataset. We also evaluated the performance of RaGS in a surround-view scenario, as shown in Table 3. While RaGS exhibits a slight performance drop in terms of the ODS metric, it consistently outperforms traditional fixed-grid models RCfusion and BEVFusion on the mAP metric. This underscores the advantage of utilizing Gaussian representations for the fusion of 4D radar and surround-view data in the context of 3D object detection.
Ablation Study
Overall Ablation. Table 4 presents an ablation study evaluating the contributions of FLI, IMA, and MGF, using Gaussianformer (Huang et al. 2024) with 4D radar as our baseline. FLI significantly aids Gaussian initialization through depth back-projection and foreground pixel selection. IMA effectively combines multimodal information, as shown in Fig. 4, progressively updating Gaussian attributes to reconstruct sparse regions of interest efficiently. Finally, MGF enhances performance by fusing multi-level features from Gaussian rasterization, enabling multi-layered perception and improving overall results.
Ablation on FLI. As shown in Table 5, FLI consists of three components: frustum-based initialization, integration of 4D radar, and depth back-projection. The frustum-based initialization helps avoid resource wastage from out-of-view objects, while the 4D radar and depth back-projection provide prior knowledge, reducing the impact of unstable segmentation and depth estimation. Additionally, the coarse but structured initialization from FLI is further refined through IMA for eliminating instability.
Ablation on IMA. IMA aggregates features from both image and radar modalities, and subsequently updates the gaussian positions within the frustum. Table 6 presents the contributions of these components. The integration of radar sparse pillars shows a relatively modest improvement, as feature aggregation has already been performed during the FLI phase and at the cross modal fuser.
Ablation on Hyper-parameters and Losses. We conduct an ablation study on the fusion level $L$ in MGF, the aggregation number $M$ in IMA, and the anchor number $N$. Both mAP and GFLOPS are reported. To balance performance and computational cost, we select 2-level MGF fusion, 3 layers for IMA, and 12,800 Gaussian anchors. See Table S1 and S2 in the supplementary material for further details.
Models $\text{mAP}_{\text{VoD}}$ $\text{mAP}_{\text{TJ4D}}$ Param. FPS GFLOPs
RCFusion 49.65 33.85 55.3M 9.0 680.8
LXL 56.31 36.32 64.0M 6.1 1169.2
SGDet3D 59.75 41.82 73.6M 9.2 1309.3
RaGS 61.86 41.72 56.5M 10.5 781.0
Table 7: Comparison with state-of-the-art models.
Metrics Extra. FLI IMA MGF All
Param. (M) 38.4 1.9 5.9 10.3 56.5
FPS 48.3 376.8 15.9 109.1 10.5
GFLOPS 508.5 4.2 162.8 98.3 781.0
Table 8: Analysis of efficiency across RaGS components.
Comparison with State-of-the-art Models. We compared RaGS with existing 4D radar and camera fusion models on mAP, FPS, and GFLOPS. As shown in Table 7, RaGS outperforms with fewer parameters and lower GFLOPS. Unlike BEV-grid methods, RaGS uses iterative refinement to focus on sparse objects while maintaining scene understanding. SGDet3D uses 160$\times$160$\times$8 fixed sampling points, about 16 times more than our 12,800. Table 8 presents efficiency analysis of each component within the RaGS Framework. The Gaussian in IMA is resource-intensive due to dynamic object focus, and future work will improve efficiency.
Conclusions
In this work, we propose RaGS, the first framework to leverage 3D Gaussian Splatting for fusing 4D radar and monocular images in 3D object detection. RaGS models the fused multi-modal features as continuous 3D Gaussians, enabling dynamically object attention while preserving comprehensive scene perception. The cascaded architecture, consisting of Frustum-based Localization Initiation (FLI), Iterative Multimodal Aggregation (IMA), and Multi-level Gaussian Fusion (MGF), progressively constructs and refines the Gaussian field through frustum-based localization, cross-modal aggregation, and multi-level fusion. Experiments on public benchmarks show state-of-the-art performance.
Limitations. RaGS currently works on single-frame data. Ongoing research aims to incorporate sequential modeling and utilize 4D radar for more explicit object attention.